OpenStudy (anonymous):

Simplify the expression [12+(6-3)^2]*10

OpenStudy (anonymous):

Rule 1: First perform any calculations inside parentheses. Rule 2: Next perform all multiplications and divisions, working from left to right. Rule 3: Lastly, perform all additions and subtractions, working from left to right.

OpenStudy (anonymous):

\[[12+(6-3)^2]*10 = [12+(3)^2]*10 = [12+(3*3)]*10 = [12+9]*10 = 21*10 = 210\]